OpenStudy (anonymous):

a satelite dish has a diameter of 8 feet and a focal length of 5 feet. what is the equation of the cross-section parabola?

OpenStudy (campbell_st):

Well the vertex will be at the origin, so the standard form is \[x^2 = 4ay\] a = the focal length so \[x^2 = 20y\] or \[y = \frac{1}{20} x^2\]

OpenStudy (campbell_st):

it looks like